# Département de mathématiques Université du Québec à Montréal 201, Avenue du Président-Kennedy

### v2

@@ Abstract
Fourier basis, goodness-of-fit, nuisance parameter, omnibus test, parametric hypothesis testing, probability integral transform, $U$-statistic.
@@ 1 Introduction
Within this framework, an omnibus goodness-of-fit test is designed to evaluate the overall fit of a model without targeting a specific alternative hypothesis [5]. Unlike directional tests that focus on particular types of deviations from the null hypothesis (such as skewness or kurtosis), omnibus tests are engineered to detect a broad spectrum of discrepancies. As eloquently stated by Rayner and Best [24, p. 5], “omnibus tests are intended to have moderate power against all alternatives.” This versatility makes them powerful instruments for identifying any form of lack of fit, though they might be less sensitive to specific deviations compared to more specialized tests.
@@
Our work makes several key contributions to the field of goodness-of-fit testing:
@@
    We propose a new goodness-of-fit test $T_{n}(\hat{\bm{\theta}}_{n})$ which, like the LK test, is based on $\sqrt{n}[C_{n}(\hat{\bm{\theta}}_{n}),S_{n}(\hat{\bm{\theta}}_{n})]^{\top}$ and fully exploits the covariance matrix $\Sigma(\bm{\theta})$, leading to a more efficient use of the available information and improved power. We show that the resulting quadratic-form statistic converges to a $\chi_{2}^{2}$ distribution under $\mathcal{H}_{0}$.
@@
    We demonstrate the strong power properties of the $T_{n}$ and LK tests through simulation studies. In a first analysis, we examine the normal, Student’s $t_{4}$, and exponential null distributions by generating power curves against several families of alternatives and comparing the results with classical EDF-based competitors. In a second analysis, we revisit the comprehensive simulation study of goodness-of-fit tests for the Laplace distribution conducted by Desgagné et al. [9], which evaluated $40$ competing procedures (including the LK test) against a large collection of $400$ alternative distributions. We augment this study by adding the $T_{n}$ test, which emerges as the most powerful Laplace test among the procedures considered.
@@
    For the first time, we analyze the asymptotic behavior of this class of tests under local alternatives, focusing on two representative examples. Building on recent theoretical advances by Desgagné et al. [11], these analyses provide a deeper understanding of the power properties of the proposed tests in neighborhoods of the null hypothesis. The performance of the $T_{n}$ test is then compared with that of a natural benchmark provided by Rao’s score test (also known as the Lagrange multiplier test), as well as with its asymptotically equivalent counterpart, the generalized likelihood ratio test (GLRT), for which the exact limiting distribution is also derived under the null hypothesis and under sequences of local alternatives.
The remainder of this paper is organized as follows. Section 2 details the construction of our goodness-of-fit tests, including the treatment of nuisance parameters. Section 3 presents the components required to implement the $T_{n}$ and LK tests for $11$ families of null distributions. Section 4 reports the results of our simulation studies, including the empirical size under $\mathcal{H}_{0}$, empirical power analyses for the normal, Student’s $t_{4}$, and exponential distributions, as well as a comprehensive empirical power study for the Laplace distribution. Section 5 investigates the asymptotic behavior of the $T_{n}$ test under local alternatives. Section 6 analyzes a dataset of surface temperature forecast errors from a numerical weather prediction model to illustrate the practical utility of the proposed tests. Section 7 summarizes our findings and provides future directions for research.
The appendices provide a link to the R code used to reproduce all numerical results (Appendix A), a table of constants (Appendix B), and the proofs of all results stated in the paper (Appendix C).
2 Goodness-of-fit test based on trigonometric moments
@@
All tests introduced in this paper are primarily based on the maximum likelihood (ML) estimator, as the regularity conditions required for ML estimation are generally mild and widely satisfied in practice. To illustrate the flexibility of the proposed approach, we also present the general framework based on an arbitrary estimator, and we provide two explicit implementations based on ML and on the method-of-moments (MM) estimator for the EPD distribution.
@@ 3 Trigonometric tests for specific distributions
$\mathrm{Beta}(\alpha,\beta)$ $f(x)=\frac{1}{B(\alpha,\beta)}x^{\alpha-1}(1-x)^{\beta-1}$,   $F(x)=I_{x}(\alpha,\beta)$,   $x\in(0,1)$, $\alpha>0$, $\beta>0$
@@
Notation. The following standard notations are used throughout the paper: $\phi(\cdot)$ and $\Phi(\cdot)$ denote the PDF and CDF of the standard normal distribution; $\Gamma(\cdot)$, $\psi(\cdot)$, and $\psi_{1}(\cdot)$ denote the gamma, digamma, and trigamma functions; $B(a,b)$, $B(x;a,b)$, and $I_{x}(a,b)=B(x;a,b)/B(a,b)$ denote the beta, incomplete beta, and regularized incomplete beta functions; $\gamma(a,x)$ denotes the lower incomplete gamma function, and $\Gamma_{a,b}(x)=\gamma(a,x/b)/\Gamma(a)$ denotes the CDF of a gamma distribution with shape parameter $a$ and scale parameter $b$. Finally,
$T(y,\lambda)=\frac{1}{2\pi}\int_{0}^{\lambda}\frac{e^{-y^{2}(1+t^{2})/2}}{1+t^{2}}\,\mathrm{d}t$
denotes Owen’s $T$ function.
Table 2: Score equations associated with each parameter of the 11 distribution families for ML estimation, along with MM estimators for the EPD family.
Several special cases can be derived from the 11 distribution families, in particular from the rich EPD and GG families. Log-variant, inverse-variant, and other types of equivalences can also be obtained, which greatly extend the applicability of the tests $T_{n}$ and LK. We provide below a list that is intended to be as exhaustive as possible:
@@
  • The Pareto type I distribution on $(1,\infty)$ is handled via $Y\sim\mathrm{\textbf{Pareto}}(\alpha)\iff X=\ln(Y)\sim\mathrm{exponential}(\beta=1/\alpha)$. For a general support $(\sigma,\infty)$, use $X=\ln(Y/\hat{\sigma}_{n})$ with $\hat{\sigma}_{n}=\min\{Y_{1},\ldots,Y_{n}\}$.
@@
Table 3: Matrices used to compute the covariance matrix $\Sigma(\bm{\theta})$, as defined in (4) or (5), and the normalizing scalar $V(\bm{\theta})$ defined in (8), assuming that the parameter vector $\bm{\theta}$ is unknown. If a given component $k$ of $\bm{\theta}$ is assumed known, it suffices to omit column $k$ of $G$ (and of $J$ in the MM case), as well as row and column $k$ of $I$ (or of $R$). The constants $h_{i}(\cdot)$ are listed in Table 13 in Appendix B.
The first step in computing the $T_{n}$ and LK test statistics is to estimate the unknown nuisance parameters. The score equations for each parameter of the 11 distribution families are given in Table 2, assuming that all parameters are unknown, for the purpose of maximum likelihood estimation. If one or more parameters are known, the corresponding score functions are simply omitted and the known values are used instead. In addition, method-of-moments (MM) estimators are provided for the EPD family as an alternative to maximum likelihood estimation, with the restriction that, in contrast to the ML case, the parameter $\lambda$ is always assumed to be known, as indicated by the notation $\mathrm{EPD}_{\lambda}(\mu,\sigma)$.
Next, the covariance matrix $\Sigma(\bm{\theta})$ for the $T_{n}$ test statistic must be computed, as defined in (4) or (5), as well as the normalizing scalar $V(\bm{\theta})$ for the LK test, defined in (8). To this end, Table 3 reports the matrices $G(\bm{\theta})$ and $I(\bm{\theta})$ for the ML-based tests, and the matrices $G(\bm{\theta})$, $J(\bm{\theta})$, and $R(\bm{\theta})$ for the MM-based case, assuming that all parameters are unknown. See Sections C.1 to C.12 of Appendix C for the derivation of all matrices reported in Table 3.
If a given component $k$ of $\bm{\theta}$ is assumed known, it suffices to omit column $k$ of $G$ (and of $J$ in the MM case), as well as row and column $k$ of $I$ (or of $R$). This yields a different covariance matrix $\Sigma(\bm{\theta})$ for each configuration of known and unknown parameters, and consequently a different test. For most distribution families, the resulting matrix $\Sigma(\bm{\theta})$ depends only on a subset of the parameters. Therefore, to simplify the expressions, the remaining parameters are set to standard values. For example, in the EPD case, we set $\mu=0$ and $\sigma=1$ in $G(\lambda,0,1)$ and $I(\lambda,0,1)$, reflecting the fact that $\Sigma(\bm{\theta})$ depends solely on $\lambda$.
@@
$I(\bm{\theta})=[I(2,0,1)]_{2:3,2:3}=\begin{bmatrix}1&0\\[2.84526pt] 0&2\end{bmatrix}$.
The constants $h_{i}(\cdot)$ are listed in Table 13 in Appendix B and are computed numerically using the R software. Using equations (4) and (8), we then obtain
@@ 4 Empirical size and power
In Section 4.1, the empirical size under $\mathcal{H}_{0}$ is investigated. We show that the chi-square $\chi_{2}^{2}$ approximation to the distributions of the $T_{n}$ and LK statistics is very accurate for small sample sizes such as $n=30$, for moderate sample sizes such as $n=100$, and asymptotically as $n\to\infty$. We conclude that critical values and $p$-values for the $T_{n}$ and LK tests can be computed with high precision without relying on Monte Carlo simulations or tabulated values.
In Section 4.2, empirical power under $\mathcal{H}_{1}$ is studied. We conduct a simulation study comparing the power of the $T_{n}$ and LK tests with classical omnibus goodness-of-fit procedures based on the probability integral transform, namely the Anderson–Darling, Cramér–von Mises, Kuiper, and Watson tests. Empirical powers are evaluated for three families of null distributions, namely the normal, Student’s $t_{4}$, and exponential distributions, using a sample size of $n=50$ and a significance level of $\alpha=0.05$. The results illustrate that the $T_{n}$ and LK tests perform very well; however, definitive conclusions would require a more extensive investigation.
In Section 4.3, we revisit the comprehensive simulation study of goodness-of-fit tests for the Laplace distribution presented by Desgagné et al. [9], which evaluated $40$ competing goodness-of-fit tests against a broad set of $400$ alternative distributions and included the MM-based LK test proposed by Langholz and Kronmal [17]. Our new MM-based Laplace test $T_{n}$ was added to this set of contenders. Given the scope and thoroughness of this study, and the fact that the $T_{n}$ test ranked first overall, we conclude that, for the specific problem of testing the Laplace distribution, the proposed $T_{n}$ test is among the most powerful omnibus goodness-of-fit tests currently available in the literature.
@@ 4.1 Empirical size under $\mathcal{H}_{0}$
In this section, we show in Tables 4 and 5 that the chi-square $\chi_{2}^{2}$ approximation to the distributions of the $T_{n}$ and LK statistics is remarkably accurate even for small sample sizes such as $n=30$, with the accuracy naturally improving as the sample size increases (e.g., $n=100$). This level of accuracy was studied across a wide range of distributions, including the normal (also covering the log-normal), logistic (also covering the log-logistic), Student’s $t$, exponential (also covering the Rayleigh and Pareto), gamma (also covering the inverse-gamma and Nakagami), and Weibull (also covering the Fréchet and Gumbel) distributions. As a result, critical values and $p$-values for the $T_{n}$ and LK tests can be computed with high precision without relying on Monte Carlo simulations or tabulated values.
@@
The empirical rejection probabilities under $\mathcal{H}_{0}$ were obtained by simulating samples of sizes $30$ and $100$ and computing the test statistics of the $T_{n}$ and LK tests. For each test, we estimated the empirical sizes at the $1\%$, $5\%$, and $10\%$ nominal levels by comparing the test statistics with the corresponding $\chi_{2}^{2}$ critical values. These Monte Carlo rejection rates, based on 100,000100{,}000 replications, are reported in Tables 4 and 5. The empirical infinite-sample sizes are also reported for the $T_{n}$ test (which equal $1\%$, $5\%$, and $10\%$ due to its $\chi_{2}^{2}$ limiting distribution) and for the LK test, the latter being computed by simulation since, under $\mathcal{H}_{0}$, $\sqrt{n}[C_{n}(\hat{\bm{\theta}}_{n}),S_{n}(\hat{\bm{\theta}}_{n})]^{\top}\rightsquigarrow\mathcal{N}_{2}(\bm{0}_{2},\Sigma(\bm{\theta}_{0}))$. We observe that although the LK statistic is not exactly asymptotically $\chi_{2}^{2}$-distributed, the chi-square approximation remains very accurate.
In Table 4, the specific values of the location–scale parameters $\mu_{0}$ and $\sigma_{0}$ (normal/log-normal, logistic/log-logistic, Student’s $t$) used to generate the data under $\mathcal{H}_{0}$ are irrelevant, since the test statistics $T_{n}$ and LK are location–scale invariant.
In Table 5, the generalized gamma distribution $\mathrm{GG}(\lambda_{0},\beta_{0},\rho_{0})$ is examined under three specific cases: (1) $\lambda_{0}=1$ known, $\beta_{0}$ unknown, and $\rho_{0}$ known, which covers the exponential, Rayleigh, and Pareto families; (2) $\lambda_{0}$ unknown, $\beta_{0}$ unknown, and $\rho_{0}$ known, which covers the gamma, inverse-gamma, and Nakagami families; and (3) $\lambda_{0}=1$ known, $\beta_{0}$ unknown, and $\rho_{0}$ unknown, which covers the Weibull, Fréchet, and Gumbel families. The particular choices of the scale and power parameters $\beta_{0}$ and $\rho_{0}$ used in the simulations under $\mathcal{H}_{0}$ are likewise irrelevant, because the statistics $T_{n}$ and LK are scale- and power-invariant for the GG test.

### v3

@@ Abstract
Fourier basis, goodness-of-fit, nuisance parameter, omnibus test, hypothesis testing, probability integral transform, $U$-statistic.
@@ 1 Introduction
Within this framework, an omnibus goodness-of-fit test is designed to evaluate the overall fit of a model without targeting a specific alternative hypothesis [5]. Unlike directional tests that focus on particular types of deviations from the null hypothesis (such as skewness or kurtosis), omnibus tests are engineered to detect a broad spectrum of discrepancies. As stated by Rayner and Best [24, p. 5], “omnibus tests are intended to have moderate power against all alternatives.” This versatility makes them powerful instruments for identifying any form of lack of fit, though they might be less sensitive to specific deviations compared to more specialized tests.
@@
Our work makes several contributions to the field of goodness-of-fit testing:
@@
    We propose a new goodness-of-fit test $T_{n}(\hat{\bm{\theta}}_{n})$ which, like the LK test, is based on $\sqrt{n}[C_{n}(\hat{\bm{\theta}}_{n}),S_{n}(\hat{\bm{\theta}}_{n})]^{\top}$ but fully exploits the covariance matrix $\Sigma(\bm{\theta})$, leading to a more efficient use of the available information and improved power. We show that the resulting quadratic-form statistic converges to a $\chi_{2}^{2}$ distribution under $\mathcal{H}_{0}$.
@@
    We demonstrate the strong power properties of the $T_{n}$ and LK tests through simulation studies. In a first analysis, we examine the normal, Student’s $t_{2}$, and exponential null distributions by generating power curves against several families of alternatives and comparing the results with classical EDF-based competitors. In a second analysis, we revisit the comprehensive simulation study of goodness-of-fit tests for the Laplace distribution conducted by Desgagné et al. [9], which evaluated $40$ competing procedures (including the LK test) against a large collection of $400$ alternative distributions. We augment this study by adding the $T_{n}$ test, which emerges as the most powerful Laplace test among the procedures considered.
@@
    We analyze the asymptotic behavior of this class of tests under local alternatives, focusing on two representative examples. Building on recent theoretical advances by Desgagné et al. [11], these analyses provide a deeper understanding of the power properties of the proposed tests in neighborhoods of the null hypothesis. The performance of the $T_{n}$ test is then compared with that of a natural benchmark provided by Rao’s score test (also known as the Lagrange multiplier test), as well as with its asymptotically equivalent counterpart, the generalized likelihood ratio test (GLRT), for which the exact limiting distribution is also derived under the null hypothesis and under sequences of local alternatives.
The remainder of this paper is organized as follows. Section 2 details the construction of our goodness-of-fit tests, including the treatment of nuisance parameters. Section 3 presents the components required to implement the $T_{n}$ and LK tests for $11$ families of null distributions. Section 4 reports the results of our simulation studies, including the empirical size under $\mathcal{H}_{0}$, empirical power analyses for the normal, Student’s $t_{2}$, and exponential distributions, as well as a comprehensive empirical power study for the Laplace distribution. Section 5 investigates the asymptotic behavior of the $T_{n}$ test under local alternatives. Section 6 analyzes a dataset of surface temperature forecast errors from a numerical weather prediction model to illustrate the practical utility of the proposed tests. Section 7 summarizes our findings and provides future directions for research.
The proofs of all results stated in the paper are collected in the Omnibus goodness-of-fit tests for univariate continuous distributions based on trigonometric moments. The appendices provide a link to the R code used to reproduce all numerical results (Appendix A), as well as a table of estimators for each parameter across the 11 distribution families and a table of constants (Appendix B).
2 Goodness-of-fit tests based on trigonometric moments
@@
All tests introduced in this paper are primarily based on the maximum likelihood (ML) estimator, as the regularity conditions required for ML estimation are generally mild and widely satisfied in practice. To illustrate the flexibility of the proposed approach, we also present the general framework based on an arbitrary estimator, and we provide explicit implementations based on ML and on the method-of-moments (MM) estimator for the EPD distribution.
@@ 3 Trigonometric tests for specific distributions
$\mathrm{Beta}(\alpha,\beta)$ $f(x)=\frac{1}{B(\alpha,\beta)}x^{\alpha-1}(1-x)^{\beta-1}$,   $F(x)=I(x;\alpha,\beta)$,   $x\in(0,1)$, $\alpha>0$, $\beta>0$
@@
Notation. The following standard notations are used throughout the paper: $\phi(\cdot)$ and $\Phi(\cdot)$ denote the PDF and CDF of the standard normal distribution; $\Gamma(\cdot)$, $\psi(\cdot)$, and $\psi_{1}(\cdot)$ denote the gamma, digamma, and trigamma functions; $B(a,b)$, $B(x;a,b)$, and $I(x;a,b)=B(x;a,b)/B(a,b)$ denote the beta, incomplete beta, and regularized incomplete beta functions; $\gamma(a,x)$ denotes the lower incomplete gamma function, and $\Gamma_{a,b}(x)=\gamma(a,x/b)/\Gamma(a)$ denotes the CDF of a gamma distribution with shape parameter $a$ and scale parameter $b$. Finally, Owen’s $T$ function is denoted by
$T(y,\lambda)=\frac{1}{2\pi}\int_{0}^{\lambda}\frac{e^{-y^{2}(1+t^{2})/2}}{1+t^{2}}\,\mathrm{d}t.$
Many special cases can be derived from the 11 distribution families, in particular from the rich EPD and GG families. Log-variant, inverse-variant, limiting-case, and other types of equivalences can also be obtained, which greatly extend the applicability of the tests $T_{n}$ and LK. We provide below a list that is intended to be as exhaustive as possible:
@@
  • The Pareto type I distribution on a known support $(\sigma,\infty)$ is handled via $Y\sim\mathrm{\textbf{Pareto}}(\alpha,\sigma)\iff X=\ln(Y/\sigma)\sim\mathrm{exponential}(\beta=1/\alpha)$. If the lower endpoint $\sigma$ is unknown, we use the plug-in $\hat{\sigma}_{n}=\min\{Y_{1},\ldots,Y_{n}\}$ and set $X=\ln(Y/\hat{\sigma}_{n})$.
@@
  • For any completely specified continuous distribution, $\Sigma=\tfrac{1}{2}I_{2}$ and $V=1$, resulting in identical test statistics for $T_{n}$ and the LK test.
@@
Table 2: Matrices used to compute the covariance matrix $\Sigma(\bm{\theta})$, as defined in (4) or (5), and the normalizing scalar $V(\bm{\theta})$ defined in (8), assuming that the parameter vector $\bm{\theta}$ is unknown. If a given component $k$ of $\bm{\theta}$ is assumed known, it suffices to omit column $k$ of $G$ (and of $J$ in the MM case), as well as row and column $k$ of II (or of $R$). The constants $h_{i}(\cdot)$ are listed in Table 13 in Appendix B.
The first step in computing the $T_{n}$ and LK test statistics is to estimate the unknown nuisance parameters. The score equations or, where available, explicit estimators for each parameter of the 11 distribution families are given in Table 12 in Appendix B, assuming that all parameters are unknown, for the purpose of maximum likelihood estimation. If one or more parameters are known, the corresponding score functions or estimators are simply omitted and the known values are used instead. In addition, method-of-moments (MM) estimators are provided for the EPD family as an alternative to maximum likelihood estimation, with the restriction that, in contrast to the ML case, the parameter $\lambda$ is always assumed to be known, as indicated by the notation $\mathrm{EPD}_{\lambda}(\mu,\sigma)$.
Next, the covariance matrix $\Sigma(\bm{\theta})$ for the $T_{n}$ test statistic must be computed, as defined in (4) or (5), as well as the normalizing scalar $V(\bm{\theta})$ for the LK test, defined in (8). To this end, Table 2 reports the matrices $G(\bm{\theta})$ and $I(\bm{\theta})$ for the ML-based tests, and the matrices $G(\bm{\theta})$, $J(\bm{\theta})$, and $R(\bm{\theta})$ for the MM-based case, assuming that all parameters are unknown. See Sections S1 to S12 of the Omnibus goodness-of-fit tests for univariate continuous distributions based on trigonometric moments for the derivation of all matrices reported in Table 2.
If a given component $k$ of $\bm{\theta}$ is assumed known, it suffices to omit column $k$ of $G$ (and of $J$ in the MM case), as well as row and column $k$ of II (or of $R$). This yields a different covariance matrix $\Sigma(\bm{\theta})$ for each configuration of known and unknown parameters, and consequently a different test. For most distribution families, the resulting matrix $\Sigma(\bm{\theta})$ depends only on a subset of the parameters. Therefore, to simplify the expressions, the remaining parameters are set to standard values. For example, in the EPD case, we set $\mu=0$ and $\sigma=1$ in $G(\lambda,0,1)$ and $I(\lambda,0,1)$, reflecting the fact that $\Sigma(\bm{\theta})$ depends solely on $\lambda$.
@@
and $I(\bm{\theta})=[I(2,0,1)]_{2:3,2:3}=\mathrm{diag}(1,2)$. The constants $h_{i}(\cdot)$ are listed in Table 13 in Appendix B and are computed numerically using the R software (for example, using the function cubature::adaptIntegrate). Using equations (4) and (8), we then obtain
@@ 4.1 Empirical size under $\mathcal{H}_{0}$
In this section, we show in Tables 3 and 4 that the chi-square $\chi_{2}^{2}$ approximation to the distributions of the $T_{n}$ and LK statistics is remarkably accurate even for small sample sizes such as $n=30$, with the accuracy naturally improving as the sample size increases (e.g., $n=100$). This level of accuracy was studied across a wide range of distributions, including the normal (also covering the log-normal), logistic (also covering the log-logistic), Student’s $t$, exponential (also covering the Rayleigh and Pareto), gamma (also covering the inverse-gamma and Nakagami), and Weibull (also covering the Fréchet and Gumbel) distributions. As a result, critical values and $p$-values for the $T_{n}$ and LK tests can be computed with high precision without relying on Monte Carlo simulations or tabulated values.
@@
The empirical rejection probabilities under $\mathcal{H}_{0}$ were obtained by simulating samples of sizes $30$ and $100$ and computing the test statistics of the $T_{n}$ and LK tests. For each test, we estimated the empirical sizes at the $1\%$, $5\%$, and $10\%$ nominal levels by comparing the test statistics with the corresponding $\chi_{2}^{2}$ critical values. These Monte Carlo rejection rates, based on 100,000100{,}000 replications, are reported in Tables 3 and 4. The empirical infinite-sample sizes are also reported for the $T_{n}$ test (which equal $1\%$, $5\%$, and $10\%$ due to its $\chi_{2}^{2}$ limiting distribution) and for the LK test, the latter being computed by simulation since, under $\mathcal{H}_{0}$, $\sqrt{n}[C_{n}(\hat{\bm{\theta}}_{n}),S_{n}(\hat{\bm{\theta}}_{n})]^{\top}\rightsquigarrow\mathcal{N}_{2}(\bm{0}_{2},\Sigma(\bm{\theta}_{0}))$. We observe that although the LK statistic is not exactly asymptotically $\chi_{2}^{2}$-distributed, the chi-square approximation remains very accurate.
In Table 3, the specific values of the location–scale parameters $\mu_{0}$ and $\sigma_{0}$ (normal/log-normal, logistic/log-logistic, Student’s $t$) used to generate the data under $\mathcal{H}_{0}$ are irrelevant, since the test statistics $T_{n}$ and LK are location–scale invariant. In Table 4, the generalized gamma distribution $\mathrm{GG}(\lambda_{0},\beta_{0},\rho_{0})$ is examined under three specific cases: (1) $\lambda_{0}=1$ known, $\beta_{0}$ unknown, and $\rho_{0}$ known, which covers the exponential, Rayleigh, and Pareto families; (2) $\lambda_{0}$ unknown, $\beta_{0}$ unknown, and $\rho_{0}$ known, which covers the gamma, inverse-gamma, and Nakagami families; and (3) $\lambda_{0}=1$ known, $\beta_{0}$ unknown, and $\rho_{0}$ unknown, which covers the Weibull, Fréchet, and Gumbel families. The particular choices of the scale and power parameters $\beta_{0}$ and $\rho_{0}$ used in the simulations under $\mathcal{H}_{0}$ are likewise irrelevant, because the statistics $T_{n}$ and LK are scale- and power-invariant for the GG test.